# Sim-to-real Transfer of Deep Reinforcement Learning Agents for Online Coverage Path Planning

### v1

@@ 4.2 Smoothening the Sim-to-Real Gap
Finally, to account for higher-order Markovian dynamics, we include information from previous time steps in the observation space. While the common approach is to stack several previous observations [11, 21], it is less feasible in our setting for multiple reasons. (1) Since the pose is embedded in the egocentric maps, rapid rotations significantly alter the observed state. Thus, it may be difficult to learn the dynamics. (2) It significantly increases the model size and processing time, which are critical aspects for real-time applications. (3) It severely limits the replay buffer size, as the map observations are fairly large. Instead, we use a history of the previous actions [11, 21], which avoids the listed problems, and should be sufficient for estimating e.g. the velocity of the robot.
4.3 Optimal Strategy for Going Sim-to-Real
@@
  2. 2.
    How does the higher-order model trained solely in simulation work on the real robot and in comparison to 1.?
  3. 3.
    How does the model with fine-tuning on the real robot perform in comparison to 2. and dependent on the number of training steps?
The working hypothesis is that 3. performs better than 2. and 1. with sufficiently small time steps approximates 2..
5 Experiments
@@ 5.1 Implementation Details
Experimental setup. The training on the physical robot is conducted using a Husqvarna Research Platform (HRP) (MIT software licence) [16], equipped with an Nvidia Jetson AGX Orin development kit. The training algorithms, based on the stable-baselines3 implementation (MIT license) [27], are executed on the Jetson, with control signals sent to the HRP. The experimental environment is a $12\times 12$ meter indoor research arena. The agent receives its position from a high-precision Qualisys motion capture system [26], which comprises 12 Oqus 700+ and 8 Arqus A12 cameras.
@@
Simulation training details. We follow the training setup by Jonnarth et al. [18], and train the CPP policy in a simulated 2D environment. Specifically, we use the same environment parameters, map geometries, and curriculum learning. We utilize soft actor-critic learning [13], and train for $2-8$M iterations with learning rate $2\cdot 10^{-5}$, batch size $256$, replay buffer size $5\cdot 10^{5}$, and discount factor $\gamma=0.99$. We use a $360^{\circ}$ simulated lidar field-of-view, $24$ lidar rays, $3.5$ m lidar range, $0.15$ m coverage radius, $0.26$ m/s maximum linear velocity, $1$ rad/s maximum angular velocity, and $0.5$ s step size. In addition, to improve the fidelity of the simulator we include a $0.5$ m/s2 maximum linear acceleration, $2$ rad/s2 maximum angular acceleration, and $50$ ms action delay. Finally, we include the $10$ latest actions in the observation. The training time was between $25$ to $100$ hours on a T4 GPU and a 6226R CPU.
@@
Real-world training details. For the fine-tuning we lower the learning rate to $10^{-5}$. We keep the same step size of $0.5$ s, which allows us to perform $4$ model updates for each environment step. In order not to give a disproportionately large weight to the first environment steps, we perform pure data collection during the first $5000$ steps, i.e. without any model updates. During real-world fine-tuning we use both fixed and randomized training maps that fit within the research arena, and set the goal coverage to $99\%$.
Evaluation. To evaluate the performance of the various RL policies, we measure the times $T_{90}$ and $T_{99}$ to reach $90\%$ and $99\%$ coverage respectively. We perform the evaluation on eight evaluation maps not seen during training, see Appendix A.
5.2 Comparison of First-order CPP Policies
@@
We first investigate how well policies that assume a first-order Markov process transfer to our semi-virtual environment. The observation space does not include any previous actions, and the policy is trained for $8$M steps in a simulated environment without a limit on the linear and angular accelerations, and does not include action delay. In Table 2, we evaluate the policy in simulation and in the real environment both before and after fine-tuning. Since this policy does not depend on the previous action sequence, we run the inference as fast as possible, i.e. we do not wait for state selection. The inference time step was measured to be 15 ms. We observe faster coverage times in the semi-virtual environment compared to the simulation. This is likely a result of the faster inference speed. The agent can update its action at a higher frequency, and can thus navigate more efficiently and produce a smoother pattern.
@@
However, after fine-tuning, we find that the performance degrades. Our hypothesis is that as the observation includes no information about the velocity, the optimal action cannot be deduced. If the robot completely changes the direction of rotation, it takes some time before the new rotation takes effect. Thus, the observation may not change much while the optimal action does. It may take a long time for the model to adapt to this, and could degrade performance initially. Running the inference at a high frequency partly circumvents this problem, explaining the high performance of the model trained solely in simulation.
5.3 Comparison of Higher-order CPP Policies
Next, we train a new model for $2$M steps in simulation, where we include the past 10 actions in the observation, and limit the maximum linear and angular accelerations. Since the 10 actions correspond to specific points back in time, we keep the inference time step the same as during training. This model can learn the higher-order Markovian dynamics, and in Table 4 we see that it transfers well when directly deployed from simulation, even under this lower inference frequency. However, also in this case does the performance degrade after fine-tuning.
To further investigate this issue, we replicate the previous experiment, and include the measured real-world overhead as an action delay in simulation to better align the simulation with reality. To avoid potential instability issues, we also freeze the feature extractors, and further lower the learning rate to $5\cdot 10^{-6}$. In Table 4 we still observe that the performance degrades after fine-tuning.
To better understand this result, we read out the learned temperature parameter in SAC [13], and find that it keeps increasing during fine-tuning. This shows that the entropy increases during training, and that the agent prioritises exploration over exploitation. In the learned paths we see that the fine-tuned agents struggle with covering the final parts of the environments. Thus, we believe that vastly longer fine-tuning is needed to learn the distribution shift to the real world, for these types of CPP policies.
Table 2: Experiment 1: 8M simulation training steps, no action observation, no inertia, no action delay. Time in minutes for reaching $90\%$ and $99\%$ coverage. “Steps”: Fine-tuning steps.
Map 1 Map 2 Map 3 Map 4 Map 5 Map 6 Map 7 Map 8 Total
Steps $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$
Simulation environment, 500 ms time step
0 6.1 8.0 5.2 7.8 5.7 9.1 5.9 9.6 5.2 8.1 6.4 9.7 8.4 11.7 8.1 11.8 51.0 75.7
Real environment, 15 ms time step
0 5.3 6.9 5.1 7.6 5.5 10.5 5.0 6.8 5.2 6.8 5.4 9.1 7.1 12.4 6.5 11.0 45.1 71.2
80k 6.4 8.4 6.1 11.0 5.6 11.6 5.7 9.0 6.2 10.4 7.0 11.8 8.3 >20 7.5 15.1 52.9 97.2
Table 3: Experiment 2: 2M simulation training steps, with action observation and inertia, and no action delay. Time in minutes for reaching $90\%$ and $99\%$ coverage. “Steps”: Fine-tuning steps.
Map 1 Map 2 Total
Steps $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$
Simulation environment, 500 ms time step
0 5.8 7.9 5.7 8.4 11.5 16.3
Real environment, 500 ms time step
0 6.2 8.0 6.4 9.2 12.6 17.2
5k 7.0 14.0 6.9 10.4 13.9 24.4
10k 7.8 12.0 7.4 11.1 15.2 23.1
15k 6.8 13.5 6.7 11.1 13.5 24.6
Table 4: Experiment 3: 2M simulation training steps, with action observation, inertia, and action delay, frozen feature extractor, and lower learning rate. Time in minutes for reaching $90\%$ and $99\%$ coverage. “Steps”: Fine-tuning steps.
Map 1 Map 2 Total
Steps $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$
Simulation environment, 500 ms time step
0 5.6 8.6 5.7 8.7 11.3 17.3
Real environment, 500 ms time step
0 6.8 10.5 7.9 10.7 14.7 21.2
5k 7.8 11.4 6.8 18.9 14.6 30.3
6 Conclusions
In this paper, we transfer a state-of-the-art CPP policy to a semi-virtual environment in order to reduce the sim-to-real gap. We find that simply training the model in simulation assuming a first-order Markov process, and deploying it with a high inference frequency is sufficient. In our experiments, the best coverage speed is attained in this way, while fine-tuning the model in the real setting degrades performance. A large amount of training time is likely required to to overcome the observed performance drop.
@@ Acknowledgements
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg (KAW) Foundation. The work was funded in part by the Vinnova project, human-centered autonomous regional airport, Dnr 2022-02678. The computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725, and by the Berzelius resource, provided by the KAW Foundation at the National Supercomputer Centre (NSC).
Appendix A Evaluation Maps
@@
Figure 3 shows the eight evaluation maps used to measure $T_{90}$ and $T_{99}$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Map 1 Map 2 Map 3 Map 4 Map 5 Map 6 Map 7 Map 8
Figure 3: The eight evaluation maps used in our experiments.

### v2

@@ Introduction
To account for higher-order Markovian dynamics, we include information from previous time steps in the observation space. While the common approach is to stack several previous observations (Haarnoja et al. 2019; Mnih et al. 2013), it is less feasible in our setting for multiple reasons. (1) Since the pose is embedded in the egocentric maps, rapid rotations significantly alter the observed state, making it difficult to learn the dynamics. (2) It significantly increases the model size and processing time, which are critical aspects for real-time applications. (3) It severely limits the replay buffer size, as the map observations are fairly large. Instead, we use a history of the previous actions (Haarnoja et al. 2019; Mnih et al. 2013), which is lightweight and avoids the listed problems, and should be sufficient for learning the dynamics of the robot. A benefit of using action observations instead of velocity estimates is that the estimates can be highly noisy. Furthermore, using action observations, the agent can learn to estimate the velocity if necessary.
Optimal Strategy for Going Sim-to-Real
@@
  2. (2)
    How does the higher-order model trained solely in simulation work on the real robot and in comparison to (1)?
  3. (3)
    How does the model with fine-tuning on the real robot perform in comparison to (2) and dependent on the number of training steps?
The working hypothesis is that (3) outperforms (2), and that (1) approximates (2) with sufficiently small step size.
Experiments
@@
Experimental setup. The training on the physical robot is conducted using a Husqvarna Research Platform (HRP) (MIT software licence) (Husqvarna 2017): a robotic lawnmower that we equip with an Nvidia Jetson AGX Orin development kit. The wheel dimensions in (1) are $r_{w}=12.25$ cm and $b=46.5$ cm. The training algorithms, based on the stable-baselines3 implementation (MIT license) (Raffin et al. 2021), are executed on the Jetson, with control signals sent to the HRP. The experimental environment is a $12\times 12$ meter indoor research arena. The agent receives its position from a high-precision Qualisys motion capture system (Qualisys 2023), which comprises 12 Oqus 700+ and 8 Arqus A12 cameras.
@@
Simulation training details. We follow the training setup by Jonnarth, Zhao, and Felsberg (2024), and train the CPP policy in a simulated 2D environment. Specifically, we use the same environment parameters (also listed in the appendix), map geometries, and curriculum learning. We utilize soft actor-critic learning (Haarnoja et al. 2018b), and train for $2$-$8$M iterations with learning rate $2\cdot 10^{-5}$, batch size $256$, replay buffer size $5\cdot 10^{5}$, and discount factor $\gamma=0.99$. We use a $360^{\circ}$ simulated lidar field-of-view, $24$ lidar rays, $3.5$ m lidar range, $0.15$ m coverage radius, $0.26$ m/s maximum linear velocity, and $1$ rad/s maximum angular velocity. To improve the fidelity of the simulator we include a $0.5$ m/s2 maximum linear acceleration, $2$ rad/s2 maximum angular acceleration, $50$ ms action delay, and include the $10$ latest actions in the observation. The training time ranged from $25$ to $100$ hours on a T4 GPU and a 6226R CPU.
@@
Real-world training details. For the fine-tuning, we lower the learning rate to $5\cdot 10^{-6}$ to avoid potential instability issues. In order not to give a disproportionately large weight to the first environment steps, we perform pure data collection during the first $5000$ steps, i.e. without any model updates. During real-world fine-tuning we use both fixed and randomized training maps that fit within the research arena, and set the goal coverage to $99\%$ for episode termination.
Evaluation. To evaluate the various RL policies, we measure the times $T_{90}$ and $T_{99}$ to reach $90\%$ and $99\%$ coverage, respectively. These metrics were chosen as the coverage time is usually important in CPP tasks, and common in the literature (Xu et al. 2022). To provide additional insights, we also measure path length and accumulated rotations, which are also of interest in some applications (Chen et al. 2019). We perform the evaluation on six maps not seen during training, see the appendix.
Extended analysis. We provide an extended analysis in the appendix, including videos, learned paths, collision statistics, performance on the individual evaluation maps, average speed and the learned entropy coefficient during fine-tuning, and an ablation study for the optimal training step size. The optimal training step size was $500$ ms, based on which, the state selection wait time in Table 1 was chosen. Finally, a step size of $500$ ms allows us to perform $4$ model updates for each environment step in the real setting.
Comparison of First-order CPP Policies
@@
We first investigate how well policies that assume a first-order Markov process transfer to our semi-virtual environment. In this experiment, the observation space does not include any previous actions, and the policy is trained for $8$M steps in a simulated environment without a limit on the linear and angular accelerations, and does not include any action delay. We evaluate the policy in simulation and in the real environment before and after fine-tuning. Since this policy does not depend on the previous action sequence, we can run it at any frequency. We evaluate it at the training frequency and as fast as possible, i.e. without waiting during state selection, which was measured to 15 ms per time step.
By comparing coverage times between simulation and reality in Table 2, we see that they are in the same ball park, where the real results are slightly worse at the lower frequency. This is expected due to differences in kinematics and other sources of error. When increasing the inference frequency, we observe faster coverage times, even surpassing the simulation results. This is likely the case due to the fact that the agent can update its actions faster and quickly adapt to errors. Thus, it can navigate more efficiently and produce a smoother pattern.
Env Steps $\Delta t$ $T_{90}$ $T_{99}$
Sim 0 500 6.0 9.1
Real 0 500 7.0 10.6
Real 0 15 5.4 8.3
Real 80k 15 6.3 10.9
Table 2: First-order policy. Average time in minutes for reaching $90\%$ and $99\%$ coverage on six evaluation maps. Env: Evaluation environment. Steps: Fine-tuning steps. $\Delta t$: Inference step size (ms). Note: Training step size is 500 ms.
@@
However, after fine-tuning, we find that the performance degrades. Our hypothesis is that as the observation includes no higher-order information, the optimal action cannot be deduced. For example, if the robot completely changes the direction of rotation, it takes some time before the new rotation takes effect. The observation may not change much, while the optimal action does. Thus, the performance degrades initially, which may take a long time to recover from. Running the inference at a high frequency partly circumvents this problem, explaining the high performance of the model without fine-tuning. While we did not evaluate the fine-tuned policy on a lower frequency, we expect it to perform worse as in the non fine-tuning case.
Figure 3: Average coverage times during fine-tuning on evaluation maps 1 and 2. Dashed line: No fine-tuning.
Figure 4: Average path length and accumulated full rotations during fine-tuning on evaluation maps 1 and 2.
Comparison of Higher-order CPP Policies
Next, we conduct a new experiment to evaluate how higher-order policies compare. We train a new model for $2$M steps in simulation, where we include the past 10 actions in the observation, and limit the maximum linear and angular accelerations based on realistic estimates. Furthermore, we include the measured real-world overhead as an action delay in simulation to better align the simulation with reality. Since the 10 actions correspond to specific points back in time, we keep the inference time step the same as during training.
The new model can learn higher-order Markovian dynamics, and in Table 3 we see that it transfers well when directly deployed from simulation, surpassing the first-order model for $99\%$ coverage under this lower inference frequency. Again, the real results are somewhat worse compared to the simulation results, as expected. In Figure 3, where we record the performance on a subset of the maps during fine-tuning, we observe that the performance degrades initially, but in contrast to the first-order model, the agent surpasses its original performance. $T_{99}$ degrades heavily in the early stage, while $T_{90}$ remains more or less constant. This suggests that the agent initially sacrifices long-term planning in favour of low-level control. After it has adapted to the low-level controls, then the long-term planning also improves. Comparing Tables 2 and 3, the higher-order policy surpasses the first-order policy in both $T_{90}$ and $T_{99}$ at $500$ ms after fine-tuning.
Furthermore, Figure 4 shows the path length and accumulated full rotations at $99\%$ coverage. The number of rotations were computed by summing the absolute differences in heading angle in every time step. Similar to the coverage time, both metrics degrade initially, and then improve over time, surpassing the initial performance. This shows that the agent finds a more efficient path after fine-tuning. It reduces the number of turns, which take time, while it overlaps its previous path less, thus reducing the total path length.
In total, the fine-tuning took roughly $8$ hours, which is only a small fraction of the weeks it would have taken to train from scratch. Note that, while training requires more computational power, such as a GPU, the computational requirements at inference time are much lower, and can even be done on a CPU (Jonnarth, Zhao, and Felsberg 2024).
Env Steps $\Delta t$ $T_{90}$ $T_{99}$
Sim 0 500 5.8 8.5
Real 0 500 7.3 10.3
Real 60k 500 6.7 10.2
Table 3: Higher-order policy. Average time in minutes for reaching $90\%$ and $99\%$ coverage on six evaluation maps. Env: Evaluation environment. Steps: Fine-tuning steps. $\Delta t$: Inference step size (ms). Note: Training step size is 500 ms.
Conclusions
In this paper, we transfer state-of-the-art CPP policies to a semi-virtual environment in order to reduce the sim-to-real gap. We address the three challenges of sim-to-real transfer presented in the introduction. (1) The mismatch between simulated and real dynamics is reduced by a high inference frequency and fine-tuning the model in a real setting. (2) Non-Markovian dynamics are accounted for by using action observations, and incorporating realistic accelerations and delays into the simulation. (3) Computational delays are addressed by parallelizing the environment interactions and the model updates. In our experiments, we find that training the model in simulation assuming a first-order Markov process is sufficient, as long as the inference frequency is sufficiently high. Meanwhile, a higher-order policy can be further improved through fine-tuning, and can be deployed at a lower inference frequency, which lowers the computational requirements for a deployed system.
Limitations and future work relate to accurate pose, simulated sensor data, and static environments with no moving obstacles. We discuss these further in the appendix.
@@ Acknowledgements
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg (KAW) Foundation. The computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725, and by the Berzelius resource, provided by the KAW Foundation at the National Supercomputer Centre (NSC).
Appendix
This appendix contains the evaluation maps, extended results and analysis, as well as a discussion on limitations.
Evaluation Maps
@@
Figure 5 illustrates the six evaluation maps used to measure the performance of the CPP policies.
Refer to caption Refer to caption Refer to caption
Map 1 Map 2 Map 3
Refer to caption Refer to caption Refer to caption
Map 4 Map 5 Map 6
Figure 5: The six evaluation maps used in our experiments.
Learned Paths
Figure 6 shows learned paths on four of the evaluation environments, and videos can be found online. [^1] Videos are available at: https://drive.google.com/drive/folders/1J0vpOBuRhCHOxsnAE1Vd1cZ6-qHPlvUh?usp=sharing
Optimal Training Step Size
In Table 4, we investigate the impact of the training step size in simulation, both for first-order and higher-order policies. In both cases, $500$ ms seems to be the best choice, considering $90\%$ and $99\%$ coverage times. Thus, we chose $500$ ms for the experiments in the main paper.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: Learned paths showing the starting position (red triangle) and the end position (green square).
First order Higher order
$\Delta t$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$
150 7.1 12.7 7.6 13.7
250 6.4 9.3 6.4 10.1
500 6.0 9.1 5.8 8.5
1000 6.1 9.6 6.4 9.1
Table 4: Training step size comparison ($\Delta t$ in ms). $T_{90}$ and $T_{99}$: Average time in minutes for reaching $90\%$ and $99\%$ coverage on the six evaluation maps in simulation.
Map 1 Map 2 Map 3 Map 4 Map 5 Map 6 Average
Env Steps $\Delta t$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$
Sim 0 500 6.1 8.0 5.2 7.8 5.7 9.1 5.9 9.6 5.2 8.1 8.1 11.8 6.0 9.1
Real 0 500 5.7 9.5 6.5 10.5 6.5 9.7 9.0 10.9 6.0 9.1 8.5 13.8 7.0 10.6
Real 0 15 5.3 6.9 5.1 7.6 5.5 10.5 5.0 6.8 5.2 6.8 6.5 11.0 5.4 8.3
Real 80k 15 6.4 8.4 6.1 11.0 5.6 11.6 5.7 9.0 6.2 10.4 7.5 15.1 6.3 10.9
Table 5: First-order policy. Time in minutes for reaching $90\%$ and $99\%$ coverage. Env: Evaluation environment. Steps: Fine-tuning steps. $\Delta t$: Inference step size (ms). Note: Training step size is always 500 ms.
Map 1 Map 2 Map 3 Map 4 Map 5 Map 6 Average
Env Steps $\Delta t$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$
Sim 0 500 5.6 8.6 5.7 8.7 5.2 8.2 5.3 7.6 5.2 7.3 7.7 10.3 5.8 8.5
Real 0 500 6.8 10.5 7.9 10.7 6.7 10.9 7.0 9.1 6.0 8.3 9.4 12.5 7.3 10.3
Real 60k 500 6.7 9.2 6.5 9.5 6.1 11.0 6.2 9.1 6.2 9.2 8.4 13.0 6.7 10.2
Table 6: Higher-order policy. Time in minutes for reaching $90\%$ and $99\%$ coverage. Env: Evaluation environment. Steps: Fine-tuning steps. $\Delta t$: Inference step size (ms). Note: Training step size is always 500 ms.
If the step size is small, it may take longer to train in simulation since the computation time per step is the same. More steps are needed to complete the same number of episodes. In Table 4, we used the same number of training steps, which meant that the number of episodes were lower for the smaller step sizes. Meanwhile, with a lower step size, future rewards are more strongly discounted, and thus, long-term planning might be limited. It is possible that this could be remedied by tuning hyperparameters, such as the discount factor, $\gamma$. We simply used the default value of $\gamma=0.99$ for SAC (Haarnoja et al. 2018a). Finally, a lower step size requires a higher number of good consecutive actions to a achieve a good result, which might inhibit learning.
On the other hand, if the step size is too large, the short-term control is limited, as each action is applied over a longer time period. This requires the agent to be more precise in its actions, which makes the problem harder. A step size of $500$ ms seemed to strike a good balance.
Extended Coverage Results
Tables 5 and 6 are extensions of Tables 2 and 3 in the main paper. They contain the $90\%$ and $99\%$ coverage times for each individual evaluation map, as well as the average times.
Average Speed
In Figure 7, we analyse the average speed of the agent during fine-tuning. The initial drop in performance is also evident here. In the initial fine-tuning phase, the agent attains a lower average speed, as a result of e.g. more collisions and many small turns, while it adjusts to the new environment dynamics. Over time, it learns to control the robot in a more efficient manner, allowing it to keep a higher average speed with less turns, which results in lower coverage times.
Learned Entropy
Figure 8 shows the learned entropy coefficient in the soft actor-critic framework (Haarnoja et al. 2018b). It increases initially, suggesting that the agent prioritizes exploration over exploitation during the distribution shift to the real world. After the initial peak, the entropy coefficient decreases, putting more emphasis on exploitation. However, it does not reach the same level as in simulation, i.e. at 00 fine-tuning steps. This either suggests that the real environment is more uncertain, or that the learning process has not entirely converged yet. It is likely a combination of the two.
Figure 7: Average speed during fine-tuning on evaluation maps 1 and 2. Dashed line: No fine-tuning.
Figure 8: The learned entropy coefficient in soft actor-critic (Haarnoja et al. 2018b) during fine-tuning.
Collision Frequency
The collision frequency on the six evaluation maps in the real setting was measured to be once every $6.2$ minutes, or once every $65.4$ meters, on average. However, the collisions were mainly from the side of the robot, and not head on. This is in line with the simulation results by Jonnarth, Zhao, and Felsberg (2024).
Additional Hyperparameters
In Table 7, we list additional hyperparameters used in our experiments. They are the same as used by Jonnarth, Zhao, and Felsberg (2024). Refer to the cited work for additional details on these parameters.
Parameter Description Value
Multi-scale maps
- grid resolution* (meters per pixel) $0.0375$
- map size** (number of grid cells) $32\times 32$
$m$ number of scales $4$
$s$ scale factor $4$
Reward parameters
$\lambda_{\mathrm{area}}$ maximum coverage reward $1$
$\lambda_{\mathrm{TV}}$ total variation reward scale $1$
$R_{\mathrm{coll}}$ collision penalty $-10$
$R_{\mathrm{const}}$ constant reward $-0.1$
Table 7: A list of additional hyperparameters used in our experiments. *Resolution for the finest scale, which is the same as the resolution for the full global map. **Note: this is not the size of the full environment, but rather the size of each scale in the multi-scale map representation.
Limitations
Accurate pose. The experiments are conducted with highly accurate pose estimates from an indoor positioning system, which may not be available, depending on the application. In such cases it needs to be estimated by other means, such as SLAM, which introduces higher levels of noise. Although the pose estimates in our experiments contain a low level of noise, Jonnarth, Zhao, and Felsberg (2024) show that adding high levels of localization and sensing noise does not impact performance to a significant extent. Thus, we believe that our conclusions likely extend to real environments with higher noise levels.
Sensing. The sim-to-real gap is mainly reduced with respect to robot kinematics and dynamics, while the sensor data remains simulated. This was done mainly to evaluate the sim-to-real transfer capability of the learning algorithm in a controlled manner, as stated in the main paper. Nevertheless, we believe that our conclusions likely extend to real sensors as well, for the same reason as above. Meanwhile, the sensing is omnidirectional in our experiments, which can be a strong assumption for some applications. However, the authors of the considered RL approach (Jonnarth, Zhao, and Felsberg 2024) demonstrated that it works even under non-omnidirectional sensing.
Dynamic obstacles. While dynamic obstacles are not explicitly modeled in the training framework, they can be accounted for by including, for instance, an object detector, to update the map based on detected moving obstacles. When an object is detected, the corresponding location in the map is replaced with non-free space in the obstacle map, and when the object is no longer detected, it is replaced with free space again. This was briefly tested in our experiments, with the following results. The robot did not collide with detected obstacles, but rather it planned around them. When the objects were moved away, and the robot explored the same region again, it planned its path through that region if necessary. In future work, we have to consider such experiments more systematically.